How many four digit pin codes begin and end with the same digit (ex 3213).
Answer #1

lets say stating with 0 and ending with 0

second digit could be 0/1/2/3/4/5/6/7/8/9

similarly third digit could be 0/1/2/3/4/5/6/7/8/9

So total number of cases would be = 10*10 = 100 when stating with 0 and ending with 0

Similarly total number of cases stating with 1 and ending with 1 would be 100 and so on......

So total cases would be = (100*10) = 1000

Please let me know if you have any doubts. Thanks!!